# Reweighted Spectral Partitioning Works: Bounds for Special Graph Classes

### v3

@@ Appendix C Deferred Proof of Subdivision Lemma
$\sum_{u,v\in V}\sum_{u^{\prime}\in p(u),v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}$
$=\sum_{u^{\prime},v^{\prime}\in V^{\prime}}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}-\sum_{v\in V}\sum_{u^{\prime},v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}.$
@@
$\sum_{v\in V}\sum_{u^{\prime},v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\leq\sum_{v\in V}\sum_{u^{\prime},v^{\prime}\in p(v)}y_{H}(u^{\prime})+y_{H}(v^{\prime})$
$\leq\sum_{v\in V}2|p(v)|\sum_{v^{\prime}\in p(v)}y_{H}(v^{\prime})\in\operatorname{\mathcal{O}}_{\Delta}(4^{k}\cdot\sum_{v^{\prime}\in V^{\prime}}y_{H}(v^{\prime}))=\operatorname{\mathcal{O}}_{\Delta}(4^{k}\cdot\gamma^{(1)}(H))$
We may assume without loss of generality that $\operatorname{\mathcal{O}}_{\Delta}(4^{k}\cdot\gamma^{(1)}(H))\leq 1$, since otherwise the lemma statement follows from the fact that $\gamma^{(1)}(G)\leq 1$. Hence, $\sum_{v\in V}\sum_{u^{\prime},v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}=\Theta_{\Delta}(\sum_{u^{\prime},v^{\prime}\in V^{\prime}}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2})$. Combining our bounds together, we obtain
$\sum_{v\in V}y_{G}(v)=\sum_{v\in V}2\rho\frac{\sum_{u:uv\in E}\sum_{x\in p(v)\cap\pi_{E}^{*}(uv)}y_{H}(v)}{\sum_{x\in V}|f_{G}(x)|^{2}}$
$=4\rho n\frac{\sum_{v\in V}\sum_{u:uv\in E}\sum_{x\in p(v)\cap\pi_{E}^{*}(uv)}y_{H}(v)}{\sum_{u,v\in V}|f_{G}(u)-f_{G}(v)|^{2}}\in\operatorname{\mathcal{O}}_{\Delta}\left(8^{k}\rho n\frac{\sum_{v^{\prime}\in V}y_{H}(v^{\prime})}{\sum_{u,v\in V}|f_{H}(\pi_{V}^{*}(u))-f_{H}(\pi_{V}^{*}(v))|^{2}}\right)$
$=\operatorname{\mathcal{O}}_{\Delta}\left(8^{k}\rho n\frac{\sum_{v^{\prime}\in V}y_{H}(v^{\prime})}{\sum_{u^{\prime},v^{\prime}\in V^{\prime}}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}}\right)=\operatorname{\mathcal{O}}_{\Delta}\left(8^{k}\rho\frac{n}{n^{\prime}}\sum_{v^{\prime}\in V}y_{H}(v^{\prime})\right).$
Using the bounds we have on $\rho$ and $\frac{n}{n^{\prime}}$, we obtain that $\sum_{v\in V}y_{G}(v)\in\Theta_{\Delta}\left(4^{k}\gamma^{(1)}(H)\right)$, as desired. ∎

### v4

@@ Appendix C Deferred Proof of Subdivision Lemma
$\sum_{u\neq v\in V}\sum_{u^{\prime}\in p(u),v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}$
$\sum_{u^{\prime},v^{\prime}\in V^{\prime}}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}-\sum_{v\in V}\sum_{u^{\prime},v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}.$
For any $v\in V$, any two $u^{\prime},v^{\prime}\in p(v)$, and any $w^{\prime}\in V^{\prime}\setminus p(v)$, the triangle inequality gives $|f_{H}(u^{\prime})-f_{H}(v^{\prime})|\leq|f_{H}(u^{\prime})-f_{H}(w^{\prime})|+|f_{H}(v^{\prime})-f_{H}(w^{\prime})|$, so by Jensen’s inequality on the convex function $x\mapsto x^{2}$, $\frac{1}{2}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\leq|f_{H}(u^{\prime})-f_{H}(w^{\prime})|^{2}+|f_{H}(v^{\prime})-f_{H}(w^{\prime})|^{2}$. In other words, every squared distance within $p(v)$ can be bounded above by two squared distances crossing out of $p(v)$. So long as there are at least two vertices, there are always more of the latter type (up to factors in $\Delta$), so it follows that
$\sum_{u\neq v\in V}\sum_{u^{\prime}\in p(u),v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\in\Omega_{\Delta}\left(\sum_{u^{\prime},v^{\prime}\in V^{\prime}}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\right).$
Finally,
[…]
$\in\operatorname{\mathcal{O}}_{\Delta}\left(n\rho\cdot 8^{k}\frac{\sum_{v^{\prime}\in V^{\prime}}y_{H}(v^{\prime})}{\sum_{u\neq v\in V}\sum_{u^{\prime}\in p(u),v^{\prime}\in p(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}}\right)$
$=\operatorname{\mathcal{O}}_{\Delta}\left(n\rho\cdot 8^{k}\frac{\sum_{v^{\prime}\in V^{\prime}}y_{H}(v^{\prime})}{\sum_{u^{\prime},v^{\prime}\in V^{\prime}}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}}\right)$
$=\operatorname{\mathcal{O}}_{\Delta}\left(n\rho\cdot 8^{k}\frac{\sum_{v^{\prime}\in V^{\prime}}y_{H}(v^{\prime})}{2n^{\prime}\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}}\right)$
$=\operatorname{\mathcal{O}}_{\Delta}\left(4^{k}\frac{\sum_{v^{\prime}\in V^{\prime}}y_{H}(v^{\prime})}{\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}}\right).$
Appendix D Deferred Proof of Area Inequality
In this section, we prove Lemma 5.7.
Lemma D.1 (Restatement of Lemma 5.7).
For a sphere cap $C$ on the $d$-dimensional unit sphere embedded in $\mathbb{R}^{d+1}$, let $f$ be its centre and let $r$ be the maximum Euclidean distance in $\mathbb{R}^{d+1}$ from $f$ to a point in $C$. Assume $C$ contains at most half the surface area of the sphere. Then, $d$-dimensional surface area of the sphere cap has $\text{area}(C)\geq\frac{A_{d}r^{d}}{4^{d}}$.
The proof will use three important well-understood functions: The Gamma function $\Gamma$, the Beta function $B$, and the regularized incomplete Beta function $I$. We will not require the explicit representation of the Gamma function, but we will use Gautschi’s inequality:
Proposition D.2 (Gautschi’s inequality [Gau59, DLMF 5.6.4]).
For $x\geq 0,s\in(0,1)$, it holds that
$x^{1-s}<\frac{\Gamma(x+1)}{\Gamma(x+s)}<(x+1)^{1-s}.$
We will use the explicit forms of the Beta and regularized incomplete Beta functions:
Definition D.3.
For $z_{1},z_{2},x>0$, the incomplete Beta function is defined as $B(x;z_{1},z_{2}):=\int_{0}^{x}t^{z_{1}-1}(1-t)^{z_{2}-1}dt$. The Beta function is defined as $B(z_{1},z_{2}):=B(1;z_{1},z_{2})$, and it is also known to be equal to $\frac{\Gamma(z_{1})\Gamma(z_{2})}{\Gamma(z_{1}+z_{2})}$. The regularized incomplete Beta function is defined as $I_{x}(z_{1},z_{2}):=\frac{B(x;z_{1},z_{2})}{B(z_{1},z_{2})}$.
We also have the following important result for our purposes:
Proposition D.4 ([Li11, Equation (1)]).
For a sphere cap $C$ with height $h\leq 1$ of the $d$-dimensional unit sphere, $\text{area}(C)=\frac{1}{2}A_{d}\cdot I_{2h-h^{2}}\left(\frac{d}{2},\frac{1}{2}\right)$.
Note that the convention for dimension differs in the reference.
Proof of Lemma 5.7.
Let $h=\frac{r^{2}}{2}$ be the height of the sphere cap (see Figure 4). Then,
$B\left(2h-h^{2};\frac{d}{2},\frac{1}{2}\right)=\int_{0}^{2h-h^{2}}t^{\frac{d}{2}-1}(1-t)^{-\frac{1}{2}}dt\geq\int_{0}^{2h-h^{2}}t^{\frac{d}{2}-1}dt=\frac{2(2h-h^{2})^{\frac{d}{2}}}{d}.$
@@ Appendix D Deferred Proof of Area Inequality
$B\left(\frac{d}{2},\frac{1}{2}\right)=\frac{\Gamma\left(\frac{d}{2}\right)\cdot\Gamma\left(\frac{1}{2}\right)}{\Gamma\left(\frac{d}{2}+\frac{1}{2}\right)}=\frac{\Gamma\left(\frac{d}{2}\right)\cdot\sqrt{\pi}}{\Gamma\left(\frac{d}{2}+\frac{1}{2}\right)}<\sqrt{\pi}\left(\frac{d}{2}-\frac{1}{2}\right)^{-\frac{1}{2}},$
where the last inequality follows from Gautschi’s inequality.
Now, $2h-h^{2}=h(2-h)\geq h$ (since $h\leq 1$), so
$I_{2h-h^{2}}\left(\frac{d}{2},\frac{1}{2}\right)\geq\frac{\frac{2(2h-h^{2})^{\frac{d}{2}}}{d}}{\sqrt{\pi}\left(\frac{d}{2}-\frac{1}{2}\right)^{-\frac{1}{2}}}\geq\frac{2h^{\frac{d}{2}}}{\sqrt{\pi}d}\left(\frac{d}{2}-\frac{1}{2}\right)^{\frac{1}{2}}=\frac{2h^{\frac{d}{2}}}{\sqrt{2\pi}}\frac{\sqrt{d-1}}{d}.$
Finally,
$\text{area}(C)=\frac{1}{2}A_{d}\cdot I_{2h-h^{2}}\left(\frac{d}{2},\frac{1}{2}\right)\geq\frac{1}{2}A_{d}\cdot\frac{2h^{\frac{d}{2}}}{\sqrt{2\pi}}\frac{\sqrt{d-1}}{d}=A_{d}\cdot\frac{r^{d}}{2^{\frac{d}{2}}\sqrt{2\pi}}\frac{\sqrt{d-1}}{d}\geq A_{d}\cdot\frac{r^{d}}{4^{d}},$
where the last inequality holds since $d\geq 2$. ∎